The temperature coefficient of copper is 0.00427/°C. The resistance of the coil is 8.82 ohms

at 24°.C. Determine its resistance at 0°.C and 42°.C.


1
Expert's answer
2022-05-23T10:17:54-0400
"R(t_2)=R(t_1)(1+\alpha (t_2-t_1))"

"R(0)=R(24)(1+\alpha (0-24))\\
=8.82*(1-0.00427*24)=7.92\:\Omega"

"R(42)=R(24)(1+\alpha (42-24))\\
=8.82*(1+0.00427*18)=9.50\:\Omega"